# Spectral Super-Resolution Neural Operator with Atmospheric Radiative Transfer Prior

### v1

@@ III-B Guidance Matrix Projection
$\text{SAM}(Y,Z)=\frac{1}{N}\sum_{n}\arccos\left(\frac{Y_{\cdot n}^{\top}Z_{\cdot n}}{\|Y_{\cdot n}\|\|Z_{\cdot n}\|}\right),$ (10)
where $\|\cdot\|$ is the $L2$ norm, $Y_{\cdot n}$ denote the $n$th column vector of $Y$, that is, the discretized spectral curve of $Y$ at pixel $n$. Equivalently, it suffices to solve
$\max$ $\quad\frac{Y_{\cdot n}^{\top}Z_{\cdot n}}{\|Y_{\cdot n}\|\|Z_{\cdot n}\|}$ (11)
s.t. $\quad SY_{\cdot n}=X_{\cdot n},$
for n=1,2,,Nn=1,2,...,N. Therefore, our goal is to find an appropriate spectral estimate $Y_{\cdot n}$ within the affine space $\mathcal{A}_{n}(S,X_{\cdot n})=\{\boldsymbol{v}\in\mathbb{R}^{C}:S\boldsymbol{v}=X_{\cdot n}\}$ that minimizes the spectral angle (or equivalently maximizes the cosine similarity) with the prior $Z_{\cdot n}$, for each pixel $n=1,2,\dots,N$.
However, the non-trivial solution to problem (11) may not exist. To address this, we impose the following assumption.
@@
For all n=1,2,,Nn=1,2,...,N, the SRF matrix $S\in\mathbb{R}^{M\times C}$, prior $Z_{\cdot n}\in\mathbb{R}^{C}$, and MSI pixel value $X_{\cdot n}\in\mathbb{R}^{M}$ satisfy:
@@
    $\alpha_{n}\triangleq X_{\cdot n}^{\top}(SS^{\top})^{-1}SZ_{\cdot n}>0$;
@@
    $\gamma_{n}\triangleq X_{\cdot n}^{\top}(SS^{\top})^{-1}X_{\cdot n}>0$.
Recall that in Section III-A we assume $S$ has full row rank, ensuring that $(SS^{\top})^{-1}$ exists. Built on the assumption and inspired by recent works on range-nullspace decomposition [4, 6], we establish the following theorem, which provides the solution to (11).
@@
$Y_{\cdot n}^{*}=S^{\top}(SS^{\top})^{-1}X_{\cdot n}+\left(I-S^{\top}(SS^{\top})^{-1}S\right)\frac{\gamma_{n}}{\alpha_{n}}Z_{\cdot n}.$ (12)
@@
Let $S^{\dagger}=S^{\top}(SS^{\top})^{-1}$ denote the Moore-Penrose inverse of $S$. To facilitate the derivation, we first decompose $Y_{\cdot n}$ and $Z_{\cdot n}$ into components lying in the range space $\mathcal{R}(S^{\dagger})=\{S^{\top}(SS^{\top})^{-1}\boldsymbol{u}:\boldsymbol{u}\in\mathbb{R}^{M}\}$ and the null space $\mathcal{N}(S)=\{\boldsymbol{v}\in\mathbb{R}^{C}:S\boldsymbol{v}=\mathbf{0}\}$, i.e., the range-null space decomposition:
$Y_{\cdot n}$ $=S^{\top}(SS^{\top})^{-1}X_{\cdot n}+\left(I-S^{\top}(SS^{\top})^{-1}S\right)Y_{\cdot n}$
$\triangleq Y_{.n,\mathcal{R}}+Y_{.n,\mathcal{N}},$ (13)
@@
In this decomposition, $Y_{\cdot n,\mathcal{N}}$ is the only optimization variable, whereas $Z_{\cdot n,\mathcal{R}}$, $Z_{\cdot n,\mathcal{N}}$, and $Y_{\cdot n,\mathcal{R}}$ are fixed and known. By the orthogonality between the range space $\mathcal{R}(S^{\dagger})$ and the null space $\mathcal{N}(S)$, the objective function of problem (11) can be expressed as
$\frac{Y_{\cdot n}^{\top}Z_{\cdot n}}{\|Y_{\cdot n}\|\|Z_{\cdot n}\|}$ $=\frac{1}{\|Z_{\cdot n}\|}\frac{(Y_{\cdot n,\mathcal{R}}+Y_{\cdot n,\mathcal{N}})^{\top}(Z_{\cdot n,\mathcal{R}}+Z_{\cdot n,\mathcal{N}})}{\|Y_{\cdot n,\mathcal{R}}+Y_{\cdot n,\mathcal{N}}\|}$
$=\frac{1}{\|Z_{\cdot n}\|}\frac{Y_{\cdot n,\mathcal{R}}^{\top}Z_{\cdot n,\mathcal{R}}+Y_{\cdot n,\mathcal{N}}^{\top}Z_{\cdot n,\mathcal{N}}}{\sqrt{\|Y_{\cdot n,\mathcal{R}}\|^{2}+\|Y_{\cdot n,\mathcal{N}}\|^{2}}}.$ (15)
@@
$\max_{Y_{.n,\mathcal{N}}\in\mathcal{N}(S)}\frac{Y_{\cdot n,\mathcal{R}}^{\top}Z_{\cdot n,\mathcal{R}}+Y_{\cdot n,\mathcal{N}}^{\top}Z_{\cdot n,\mathcal{N}}}{\sqrt{\|Y_{\cdot n,\mathcal{R}}\|^{2}+\|Y_{\cdot n,\mathcal{N}}\|^{2}}}.$ (16)
The problem (16) attains its optimum if and only if $Y_{\cdot n,\mathcal{N}}$ and $Z_{\cdot n,\mathcal{N}}$ are collinear, so that their inner product in the numerator is maximized. Hence, let $Y_{\cdot n,\mathcal{N}}=\xi Z_{\cdot n,\mathcal{N}}$, $\xi\geq 0$, and the problem can be further simplified to a univariate optimization problem:
@@
$g(\xi)$ $=\frac{Y_{\cdot n,\mathcal{R}}^{\top}Z_{\cdot n,\mathcal{R}}+\|Z_{\cdot n,\mathcal{N}}\|^{2}\xi}{\sqrt{\|Y_{\cdot n,\mathcal{R}}\|^{2}+\|Z_{\cdot n,\mathcal{N}}\|^{2}\xi^{2}}}$
@@
$g^{\prime}(\xi)=\frac{\beta_{n}\gamma_{n}-\alpha_{n}\beta_{n}\xi}{\left(\beta_{n}\xi^{2}+\gamma_{n}\right)^{\frac{3}{2}}}.$ (19)
@@
$Y_{\cdot n}^{*}$ $=Y_{.n,\mathcal{R}}+\xi^{*}Z_{.n,\mathcal{N}}$
$=S^{\top}(SS^{\top})^{-1}X_{\cdot n}+\left(I-S^{\top}(SS^{\top})^{-1}S\right)\frac{\gamma_{n}}{\alpha_{n}}Z_{\cdot n}.$ (20)
Therefore, for each pixel $n=1,2,\ldots,N$, the optimum upsampled spectral curve $Y_{\cdot n}^{*}$ is obtained as given in (12). Finally, by concatenating all $Y_{\cdot n}^{*}$, we obtain the final result in matrix form and define the GMP operator as
$P_{\text{GMP}}(Z,S,X)\triangleq Y^{*}=S^{\dagger}X+P_{S}Z\Pi,$ (21)
@@
Given $X$, $S$, and $Z$, the upsampled HSI can be computed in closed form using (21), without the need for any iterative optimization or training procedure. Additionally, 1 is generally mild and commonly satisfied in practical scenarios, thus ensuring the broad applicability of the GMP method.
@@ III-C Reconstruction Neural Operator
In this section, we introduce the neural operator architecture for the reconstruction stage. In this stage, we aim to learn the continuous mapping between the upsampled HSI $\bar{Y}$ and the target HSI $Y$ via a data-driven neural network. The output $\tilde{Y}$ in (7) is expected to be more informative and more accurate than the input $\bar{Y}$, i.e., closer to the ground truth $Y$.
@@
12:return $\tilde{y}$
In general, based on the formulation of the neural operator in Section II-B, the structure of the reconstruction neural operator $R_{no}(\cdot,\boldsymbol{w})$ can be summarized as follows.
@@
    Lifting: The upsampled HSI $\bar{Y}$, together with the input coordinate $\boldsymbol{w}$, is lifted into a higher-dimensional feature space through a multilayer perceptron (MLP), formulated as
    $\tilde{Y}_{0}=\text{MLP}_{P}\left(\text{Concat}\left(\bar{Y},\boldsymbol{w}\right)\right),$ (23)
    where $\tilde{Y}_{0}$ is the first hidden representation, $\text{Concat}(\cdot)$ is the concatenation operation, and trivial reshaping and expanding operations are omitted for clarity.
@@
    $\tilde{Y}_{t+1}=\sigma\left(\mathcal{L}_{t}\left(\tilde{Y}_{t}\right)+\text{SAC}_{t}\left(\tilde{Y}_{t}\right)\right),$ (24)
@@
    $\tilde{Y}_{t+1}=\sigma\left(\mathcal{S}_{t}\left(\text{Concat}\left(\tilde{Y}_{t},\tilde{Y}_{2T_{c}+T_{r}-1-t}\right)\right)\right),$ (25)
@@
    Projection: We finally map the last hidden representation $\tilde{Y}_{2T_{c}+T_{r}}$ to the estimated HSI $\tilde{Y}$ via another MLP, that is,
    $\tilde{Y}=\text{MLP}_{Q}\left(\tilde{Y}_{2T_{c}+T_{r}}\right).$ (26)
@@ IV Experiments
In this section, we present numerical experiment results to validate our approach under various settings. We begin by introducing the datasets in Section IV-A and outlining the basic experimental setups in Section IV-B. Comparisons with several baselines are reported in Section IV-C, followed by an ablation study in Section IV-D. Finally, experiments on arbitrary-scale and zero-shot spectral reconstruction are conducted in Section IV-E to further assess the generalization capability of the model.
@@ IV-A Datasets
We collect data from the NASA Jet Propulsion Laboratory (JPL), originally acquired by the widely recognized AVIRIS (Airborne Visible/Infrared Imaging Spectrometer) sensors, and curate it into a new dataset for our study [^1] Data available at https://aviris.jpl.nasa.gov/dataportal. The dataset comprises hyperspectral images spanning 224 contiguous spectral bands, ranging from 400 to 2500 nm. Specifically, we select various typical and representative sites, each comprising multiple scenes. The scene characteristics, such as season and spatial resolution, are consistent within each site but differ significantly across different sites. The summary of sites used in our experiment is provided in Table II.
TABLE II: AVIRIS Sites Selected for the Experiments
@@ IV-B1 Basic Setup
In our experiments, we use the direct beam radiation predicted by SMARTS [11] based on the 1976 U.S. Standard Atmosphere as the ART prior, which serves as the prior $Z$ in (6). For the SRF, we utilize the database provided by [17], which includes 28 camera sensitivity functions covering a variety of types. We generate the paired RGB image by multiplying the normalized SRF randomly selected from the database and the HSI from the AVIRIS dataset we collected in Section IV-A. For hyperparameters of the neural operator, we set maximum Fourier modes $d_{\text{modes}}=16$, number of hidden channels $d=32$, and number of contracting and transformation layers $T_{c}=T_{r}=4$. Coordinate-based terms, including SRF $S_{\boldsymbol{w}}$ and ART prior $Z_{\boldsymbol{w}}$, are interpolated at wavelength $\boldsymbol{w}$ via kernel regression.
@@ IV-B2 Loss Function
$\mathcal{L}(\hat{Y},Y)=L_{1}(\hat{Y},Y)+\lambda\text{SAM}(\hat{Y},Y),$ (27)
@@ IV-B3 Metrics
IV-C Comparative Experiments
TABLE III: Comparative Results on the Overall Datasets
Method Params (M) Arbitrary-scale MRAE $\downarrow$ PSNR $\uparrow$ SAM $\downarrow$ SSIM $\uparrow$
@@ IV-C Comparative Experiments
SSRNO (Ours) 4.8 0.160 45.400 0.137 0.973
We conduct comparative experiments to evaluate the effectiveness of the proposed SSRNO against several representative methods, including UnGUN [34], INR [48], NeSR [42], and NeSSR [43]. All models are evaluated under the same training-testing protocol using the full-resolution spectral super-resolution setting.
Table III reports the overall quantitative results on the test datasets. Among all baselines, our SSRNO consistently achieves the best performance, with the lowest MRAE and SAM, the highest PSNR and SSIM, and the fewest parameters among high-performing methods. These results highlight the efficiency and effectiveness of our design. Furthermore, we also conduct per-site evaluations on six representative scenes, and the results are summarized in Table IV. SSRNO consistently delivers top or near-top performance in most scenarios, particularly excelling in complex scenes such as rural areas. This demonstrates that the model is highly generalizable and adaptable to heterogeneous landforms and spectral distributions.
We also conduct a visual comparison among all models, as shown in Figure 5, Figure 6, and Figure 7. Figure 5 demonstrates that our method is capable of faithfully restoring RGB images, whereas other methods exhibit visual artifacts or unnatural color distortions. This advantage mainly stems from the refinement step, which explicitly enforces zero reconstruction loss with GMP, ensuring accurate color consistency. The error maps in Figure 6 further highlight the superiority of our method. Additionally, we plot the predicted spectrum on six sites, which are shown in Figure 7. Our method yields more natural and physically plausible spectra that better align with the ground-truth, especially in wavelength regions affected by atmospheric absorption, whereas other approaches tend to either overfit noisy variations or underfit the true reflectance shape.
In general, both quantitative and qualitative comparisons indicate that SSRNO achieves state-of-the-art performance and more stable results across varied geographical contexts with fewer parameters.
TABLE IV: Comparative Results for Each Site
@@
SSRNO (Ours) 43.296 0.112 53.423 0.049 42.188 0.183 41.884 0.148 41.601 0.125 45.575 0.221
TABLE V: Quantitative Evaluation of Continuous (Interpolation) and Zero-shot (Extrapolation) Spectral Reconstruction
Method Continuous (Interpolation) Zero-shot (Extrapolation)
@@
SSRNO (Ours) 0.220 43.144 0.194 0.944 0.707 34.457 0.641 0.765
IV-D Ablation Study
In this section, we conduct ablation studies to further investigate the effectiveness of the proposed method. The key innovation of our method lies in the GMP, which integrates physical priors into a data-driven pipeline during the initial upsampling and enforces hard constraints during final refinement. To validate its effectiveness, we perform ablation studies on the upsampling stage and the refinement stage. For the upsampling stage, we use a zero matrix as the prior $Z$ to obtain the upsampled HSI, rather than using the ART prior, which is also equivalent to employing the Moore-Penrose inverse. To compare the effectiveness of hard and soft constraints, we directly remove the final refinement stage and add an extra error term on the loss function as the regularization to enforce low reconstruction error, which is
$\mathcal{L}_{r}(\hat{Y},Y)=\mathcal{L}(\hat{Y},Y)+\alpha L_{1}(S_{\boldsymbol{w}}\hat{Y},S_{\boldsymbol{w}}Y),$ (29)
where $\alpha$ is the trade-off parameter, and we set $\alpha=0.5$ in the ablation experiment. The results are listed in Table VI.
The results indicate that both the ART prior and the final refinement play critical roles in the overall performance. Removing the ART prior in the upsampling stage reduces PSNR and increases SAM, showing that physically informed guidance is crucial for producing an initial estimate close to the target. On the other hand, eliminating the refinement stage and replacing it with soft regularization degrades both PSNR and SAM significantly, suggesting that hard constraints in the output space are more effective in preserving spectral fidelity. These findings highlight that the GMP method plays a crucial role in ensuring accurate and physically consistent spectral reconstruction.
TABLE VI: Ablation Study of ART Prior and Refinement
ART Prior Refinement PSNR $\uparrow$ SAM $\downarrow$

### v2

@@ III-B Angular-Consistent Projection
$\text{SAM}(X,Z)=\frac{1}{N}\sum_{n}\arccos\left(\frac{X_{\cdot n}^{\top}Z_{\cdot n}}{\|X_{\cdot n}\|\|Z_{\cdot n}\|}\right),$ (10)
where $\|\cdot\|$ is the $L2$ norm, $X_{\cdot n}$ denote the $n$th column vector of $X$, that is, the discretized spectral curve of $X$ at pixel $n$. Equivalently, it suffices to solve
$\max$ $\quad\frac{X_{\cdot n}^{\top}Z_{\cdot n}}{\|X_{\cdot n}\|\|Z_{\cdot n}\|}$ (11)
s.t. $\quad Y_{\cdot n}=SX_{\cdot n},$
for n=1,2,,Nn=1,2,...,N. Therefore, our goal is to find an appropriate spectral estimate $X_{\cdot n}$ within the affine space $\mathcal{A}_{n}(S,Y_{\cdot n})=\{\boldsymbol{v}\in\mathbb{R}^{C}:S\boldsymbol{v}=Y_{\cdot n}\}$ that minimizes the spectral angle (or equivalently maximizes the cosine similarity) with the prior $Z_{\cdot n}$, for each pixel $n=1,2,\dots,N$.
However, since problem (11) is non-convex, the existence of a non-trivial solution is not guaranteed. Under the following assumption, we establish the existence of a non-trivial global optimum.
@@
For all n=1,2,,Nn=1,2,...,N, the SRF matrix $S\in\mathbb{R}^{M\times C}$, prior $Z_{\cdot n}\in\mathbb{R}^{C}$, and MSI pixel value $Y_{\cdot n}\in\mathbb{R}^{M}$ satisfy:
@@
    $\alpha_{n}\triangleq Y_{\cdot n}^{\top}(SS^{\top})^{-1}SZ_{\cdot n}>0$;
@@
    $\gamma_{n}\triangleq Y_{\cdot n}^{\top}(SS^{\top})^{-1}Y_{\cdot n}>0$.
Recall that in Section III-A we assume $S$ has full row rank, ensuring that $(SS^{\top})^{-1}$ exists. Built on the assumption and inspired by recent works on null-space decomposition [4, 7], we establish the following theorem, which provides the global optimal solution to (11).
@@
$X_{\cdot n}^{*}=S^{\top}(SS^{\top})^{-1}Y_{\cdot n}+\left(I-S^{\top}(SS^{\top})^{-1}S\right)\frac{\gamma_{n}}{\alpha_{n}}Z_{\cdot n}.$ (12)
@@
Let $S^{\dagger}=S^{\top}(SS^{\top})^{-1}$ denote the Moore-Penrose inverse of $S$. To facilitate the derivation, we first decompose $X_{\cdot n}$ and $Z_{\cdot n}$ into components lying in the range space $\mathcal{R}(S^{\dagger})=\{S^{\top}(SS^{\top})^{-1}\boldsymbol{u}:\boldsymbol{u}\in\mathbb{R}^{M}\}$ and the null space $\mathcal{N}(S)=\{\boldsymbol{v}\in\mathbb{R}^{C}:S\boldsymbol{v}=\mathbf{0}\}$, i.e., the range-null space decomposition:
$X_{\cdot n}$ $=S^{\top}(SS^{\top})^{-1}Y_{\cdot n}+\left(I-S^{\top}(SS^{\top})^{-1}S\right)X_{\cdot n}$
$\triangleq X_{.n,\mathcal{R}}+X_{.n,\mathcal{N}},$ (13)
@@
In this decomposition, $X_{\cdot n,\mathcal{N}}$ is the only optimization variable, whereas $Z_{\cdot n,\mathcal{R}}$, $Z_{\cdot n,\mathcal{N}}$, and $X_{\cdot n,\mathcal{R}}$ are fixed and known. By the orthogonality between the range space $\mathcal{R}(S^{\dagger})$ and the null space $\mathcal{N}(S)$, the objective function of problem (11) can be expressed as
$\frac{X_{\cdot n}^{\top}Z_{\cdot n}}{\|X_{\cdot n}\|\|Z_{\cdot n}\|}$ $=\frac{1}{\|Z_{\cdot n}\|}\frac{(X_{\cdot n,\mathcal{R}}+X_{\cdot n,\mathcal{N}})^{\top}(Z_{\cdot n,\mathcal{R}}+Z_{\cdot n,\mathcal{N}})}{\|X_{\cdot n,\mathcal{R}}+X_{\cdot n,\mathcal{N}}\|}$
$=\frac{1}{\|Z_{\cdot n}\|}\frac{X_{\cdot n,\mathcal{R}}^{\top}Z_{\cdot n,\mathcal{R}}+X_{\cdot n,\mathcal{N}}^{\top}Z_{\cdot n,\mathcal{N}}}{\sqrt{\|X_{\cdot n,\mathcal{R}}\|^{2}+\|X_{\cdot n,\mathcal{N}}\|^{2}}}.$ (15)
@@
$\max_{X_{.n,\mathcal{N}}\in\mathcal{N}(S)}\frac{X_{\cdot n,\mathcal{R}}^{\top}Z_{\cdot n,\mathcal{R}}+X_{\cdot n,\mathcal{N}}^{\top}Z_{\cdot n,\mathcal{N}}}{\sqrt{\|X_{\cdot n,\mathcal{R}}\|^{2}+\|X_{\cdot n,\mathcal{N}}\|^{2}}}.$ (16)
The problem (16) attains its optimum if and only if $X_{\cdot n,\mathcal{N}}$ and $Z_{\cdot n,\mathcal{N}}$ are collinear, so that their inner product in the numerator is maximized. Hence, let $X_{\cdot n,\mathcal{N}}=\xi Z_{\cdot n,\mathcal{N}}$, $\xi\geq 0$, and the problem can be further simplified to a univariate optimization problem:
@@
$g(\xi)$ $=\frac{X_{\cdot n,\mathcal{R}}^{\top}Z_{\cdot n,\mathcal{R}}+\|Z_{\cdot n,\mathcal{N}}\|^{2}\xi}{\sqrt{\|X_{\cdot n,\mathcal{R}}\|^{2}+\|Z_{\cdot n,\mathcal{N}}\|^{2}\xi^{2}}}$
@@
$g^{\prime}(\xi)=\frac{\beta_{n}\gamma_{n}-\alpha_{n}\beta_{n}\xi}{\left(\beta_{n}\xi^{2}+\gamma_{n}\right)^{3/2}}.$ (19)
@@
$X_{\cdot n}^{*}$ $=X_{.n,\mathcal{R}}+\xi^{*}Z_{.n,\mathcal{N}}$
$=S^{\top}(SS^{\top})^{-1}Y_{\cdot n}+\left(I-S^{\top}(SS^{\top})^{-1}S\right)\frac{\gamma_{n}}{\alpha_{n}}Z_{\cdot n}.$ (20)
Therefore, for each pixel $n=1,2,\ldots,N$, the optimum upsampled spectral curve $X_{\cdot n}^{*}$ is obtained as given in (12). Finally, by concatenating all $X_{\cdot n}^{*}$, we obtain the final result in matrix form and define the ACP operator as
$P_{\text{ACP}}(Z,S,Y)\triangleq X^{*}=S^{\dagger}Y+P_{S}Z\Pi,$ (21)
@@
Given $Y$, $S$, and $Z$, the upsampled HSI can be computed in closed form using (21), without the need for any iterative optimization or training procedure. Additionally, 1 is generally mild and commonly satisfied in practical scenarios, thus ensuring the broad applicability of the ACP method.
@@ III-C Reconstruction Neural Operator
In this section, we introduce the neural operator architecture for the reconstruction stage. In this stage, we aim to learn the continuous mapping between the upsampled HSI $\bar{X}$ and the target HSI $X$ via a data-driven neural network. The output $\tilde{X}$ in (7) is expected to be more informative and more accurate than the input $\bar{X}$, i.e., closer to the ground truth $X$.
@@
12:return $y$
In general, based on the formulation of the neural operator in Section II-C, the structure of the reconstruction neural operator $R_{no}(\cdot,\boldsymbol{w})$ can be summarized as follows.
@@
    Lifting: The upsampled HSI $\bar{X}$, together with the input coordinate $\boldsymbol{w}$, is lifted into a higher-dimensional feature space through a multilayer perceptron (MLP), formulated as
    $\tilde{X}_{0}=\text{MLP}_{P}\left(\text{Concat}\left(\bar{X},\boldsymbol{w}\right)\right),$ (23)
    where $\tilde{X}_{0}$ is the first hidden representation, $\text{Concat}(\cdot)$ is the concatenation operation, and trivial reshaping and expanding operations are omitted for clarity.
@@
    $\tilde{X}_{t+1}=\sigma\left(\mathcal{L}_{t}\left(\tilde{X}_{t}\right)+\text{SAC}_{t}\left(\tilde{X}_{t}\right)\right),$ (24)
@@
    $\tilde{X}_{t+1}=\sigma\left(\mathcal{S}_{t}\left(\text{Concat}\left(\tilde{X}_{t},\tilde{X}_{2T_{c}+T_{r}-1-t}\right)\right)\right),$ (25)
@@
    Projection: We finally map the last hidden representation $\tilde{X}_{2T_{c}+T_{r}}$ to the estimated HSI $\tilde{X}$ via another MLP, that is,
    $\tilde{X}=\text{MLP}_{Q}\left(\tilde{X}_{2T_{c}+T_{r}}\right).$ (26)
@@ IV Experiments
In this section, we present numerical experiment results to validate our approach under various settings. We begin by introducing the datasets in Section IV-A and outlining the basic experimental setups in Section IV-B. We first evaluate our model under the conventional discrete SSR setting, with results reported in Section IV-C. We then assess its performance in continuous spectral super-resolution and infrared spectral extrapolation scenarios in Section IV-D. Finally, an ablation study is presented in Section IV-E to analyze how and why the proposed design is effective.
@@ IV-A Datasets
We collect data from the NASA Jet Propulsion Laboratory (JPL), originally acquired by the widely recognized AVIRIS (Airborne Visible/Infrared Imaging Spectrometer) sensors, and curate it into a new dataset for our study [^1] Data available at https://aviris.jpl.nasa.gov/dataportal. The dataset comprises hyperspectral images spanning 224 contiguous spectral bands, ranging from 400 to 2500 nm. Specifically, we select various typical and representative sites, each comprising multiple scenes. The scene characteristics, such as season and spatial resolution, are consistent within each site but differ significantly across different sites. The summary of sites used in our experiment is provided in Table I.
TABLE I: AVIRIS Sites Selected for the Experiments
@@ IV-B1 Basic Setup
In our experiments, we use the direct beam radiation predicted by SMARTS [11] based on the 1976 U.S. Standard Atmosphere as the radiative spectral prior, which serves as the angular guidance $Z$ in (6). For the SRF, we utilize the database provided by [17], which includes 28 camera sensitivity functions covering a variety of types. We generate the paired RGB image by multiplying the normalized SRF randomly selected from the database and the HSI from the AVIRIS dataset we collected in Section IV-A, as shown in (4). For hyperparameters of the neural operator, we set maximum Fourier modes $d_{\text{modes}}=16$, number of hidden channels $d=32$, and number of contracting and transformation layers $T_{c}=T_{r}=4$. Coordinate-based terms, including SRF $S_{\boldsymbol{w}}$ and radiative prior $Z_{\boldsymbol{w}}$, are interpolated at wavelength $\boldsymbol{w}$ via kernel regression.
@@ IV-B2 Loss Function
$\mathcal{L}(\hat{X},X)=L_{1}(\hat{X},X)+\lambda\text{SAM}(\hat{X},X),$ (27)
@@ IV-B3 Metrics
IV-C Discrete Spectral Super-Resolution
TABLE II: Comparative Results on the Overall Datasets
Method Params (M) Continuous MRAE $\downarrow$ PSNR $\uparrow$ SAM $\downarrow$ SSIM $\uparrow$
@@ IV-C Discrete Spectral Super-Resolution
RSNO (Ours) 4.8 0.160 45.400 0.137 0.973
We conduct comparative experiments in conventional discrete SSR setting to evaluate the performance of the proposed RSNO against several representative methods, including UnGUN [35], INR [48], NeSR [45], and NeSSR [46]. All models are evaluated under the same training-testing protocol.
Table II reports the overall quantitative results on the test datasets. Among all baselines, our RSNO consistently achieves the best performance, with the lowest MRAE and SAM, the highest PSNR and SSIM, and the fewest parameters among high-performing methods. These results highlight the efficiency and effectiveness of our design. Furthermore, we also conduct per-site evaluations on six representative scenes, and the results are summarized in Table III. RSNO consistently delivers top or near-top performance in most scenarios, particularly excelling in complex scenes such as rural areas. This demonstrates that the model is highly generalizable and adaptable to heterogeneous landforms and spectral distributions.
We also conduct a visual comparison among all models, as shown in Figure 3, Figure 4, and Figure 5. Figure 3 demonstrates that our method is capable of faithfully restoring RGB images, whereas other methods exhibit visual artifacts or unnatural color distortions. This advantage mainly stems from the refinement step, which explicitly enforces zero reconstruction loss with ACP, ensuring accurate color consistency. The error maps in Figure 4 further highlight the superiority of our method. Additionally, we plot the predicted spectrum on six sites, which are shown in Figure 5. Our method yields more natural and physically plausible spectra that better align with the ground-truth, especially in wavelength regions affected by atmospheric absorption, whereas other approaches tend to either overfit noisy variations or underfit the true reflectance shape.
In general, both quantitative and qualitative comparisons indicate that RSNO achieves state-of-the-art performance and more stable results in discrete setting across varied geographical contexts with fewer parameters.
TABLE III: Comparative Results for Each Site
@@
RSNO (Ours) 43.296 0.112 53.423 0.049 42.188 0.183 41.884 0.148 41.601 0.125 45.575 0.221
TABLE IV: Quantitative Evaluation of Continuous and Extrapolative Spectral Reconstruction
Method Continuous Interpolation Infrared Extrapolation
@@
RSNO (Ours) 0.220 43.144 0.194 0.944 0.707 34.457 0.641 0.765
IV-D Continuous and Extrapolative Spectral Super-Resolution
To validate the effectiveness of our continuous modeling and evaluate the generalization and flexibility of our method, we conduct experiments on continuous spectral reconstruction and infrared spectral extrapolation.
In the continuous reconstruction setting, we train the model using only a small number of uniformly downsampled spectral bands and test it on the full spectral resolution. This simulates an interpolation task in the spectral domain, where the model needs to infer missing spectral information between known bands. In the infrared extrapolation setting, we train the model using only the visible spectrum range from 400 nm to 1000 nm, but test the model on the complete spectral range from 400 nm to 2500nm, which evaluates the ability to reconstruct the entire hyperspectral signal, including the never-seen infrared region. This corresponds to a spectral extrapolation task, which is highly challenging due to the lack of direct observation in the infrared spectral range.
In our experiments, we apply $2\times$ downsampling for continuous spectral reconstruction and restrict the spectrum to wavelengths below 1000 nm for extrapolative reconstruction. We compare our method against two other continuous methods: NeSR and NeSSR. Table IV summarizes the quantitative results. Our method achieves best performance in both settings, indicating strong generalization to unseen or sparsely-sampled spectral distributions.
The good performance of RSNO stems from two factors: the neural operator backbone ensuring resolution-invariant spectral interpolation, and the ACP method guided by radiative spectral prior ensuring radiative-structured extrapolation from visible to infrared bands. These results highlight RSNO’s robustness and generalization ability, showing that it can both interpolate missing spectra and extrapolate beyond observed bands, making it suitable for practical hyperspectral applications.
IV-E Ablation Study
TABLE V: Ablation Study of Radiative Prior and Refinement
Radiative Prior Refinement PSNR $\uparrow$ SAM $\downarrow$